Find the inverse of the matrix


1 −1 4
2 9 8
6 5 2

 using LU decomposition method.
1
Expert's answer
2020-04-21T15:32:59-0400


The LU decomposition "A=LU"  of the given matrix is:


"A=\begin{bmatrix}
 1 & -1&4 \\
 2 &9&8\\
6&5&2
\end{bmatrix}=
\begin{bmatrix}
 1 & 0&0 \\
 2 & 1&0\\
6&a&1
\end{bmatrix}\begin{bmatrix}
 1 & -1&4 \\
 0 & b&0\\
0&0&c
\end{bmatrix}"

-2+b=9 -6+11a=5 24+c=2

b=11 a=1 c=-22

"A=
\begin{bmatrix}
 1 & 0&0 \\
 2 & 1&0\\
6&1&1
\end{bmatrix}\begin{bmatrix}
 1 & -1&4 \\
 0 & 11&0\\
0&0&-22
\end{bmatrix}=LU"

Hence, the inverse of the matrix is "A^{-1}=U^{-1}L^{-1}"

Since U and L are already in upper and lower triangular form, it is easy to find their inverses using Gauss-Jordan elimination.

Inverse of L


"\begin{bmatrix}
 1 & 0&0 |1&0&0\\
 2 &1&0|0&1&0\\
6&1&1|0&0&1
\end{bmatrix}"

IIr+Ir(-2)

IIIr+Ir(-6)

"\begin{bmatrix}
 1 &0&0|\quad1&0&0\\
 0 &1&0|-2&1&0\\
0&1&1|-6&0&1
\end{bmatrix}"

IIIr+IIr(-1)

"\begin{bmatrix}
 1 &0&0 |1&0&0\\
 0 &1&\quad0 |-2&1&0\\
0&0&\quad1|-4&-1&1
\end{bmatrix}"

Inverse of U:

"\begin{bmatrix}
 1 & -1&\quad4 |1&0&0\\
 0 &11&\quad0 |0&1&0\\
0&1&-22|0&0&1
\end{bmatrix}"

IIr"\cdot(\frac{1}{11})"

"\begin{bmatrix}
 1 & -1&\quad4 |1&0&0\\
 0 &1&\quad0 |0&\frac{1}{11}&0\\
0&1&-22|0&0&1
\end{bmatrix}"

IIIr+IIr(-1)

"\begin{bmatrix}
 1 & -1&\quad4 |\frac{1}{11}&-\frac{1}{11}&\frac{2}{11}\\
 0 &1&\quad0 |0&\frac{1}{11}&0\\
0&0&-22|0&-\frac{1}{11}&1
\end{bmatrix}"


IIIr"\cdot(-\frac{1}{22})"

Ir+IIr+IIIr(-4)


"\begin{bmatrix}
 1 & 0&\quad0 |1&\frac{1}{11}&\frac{2}{11}\\
 0 &1&\quad0 |0&\frac{1}{11}&0\\
0&0&1|0&0&-\frac{1}{22}
\end{bmatrix}"

Hence,

"A^{-1}=U^{-1}L^{-1}=\\=\begin{bmatrix}
1&\frac{1}{11}&\frac{2}{11}\\
 0&\frac{1}{11}&0\\
0&0&-\frac{1}{22}
\end{bmatrix}
\begin{bmatrix}
 1&0&0\\
 -2&1&0\\
-4&-1&1
\end{bmatrix}=\\
=
\begin{bmatrix}
 \frac{1}{11}&-\frac{1}{11}&\frac{2}{11}\\
-\frac{2}{11}&\frac{1}{11}&0\\
\frac{2}{11}&\frac{1}{22}&-\frac{1}{22}
\end{bmatrix}"